OpenStudy (anonymous):

What is the limit as x approaches zero of f(x)=sin(1/x) ? Why would it be zero as opposed to undefined?

OpenStudy (anonymous):

It is undefined...

OpenStudy (cwrw238):

i think its is undefended

OpenStudy (cwrw238):

* undefined lol

OpenStudy (anonymous):

You can prove that the limit is undefined by the following method: Make the following substitution for \(\bf x\):\[\bf h=\frac{1}{x}\]Now as \(\bf x\) approaches 0, \(\bf h\) approaches \(\bf \infty\). Hence we re-write our limit as:\[\bf \lim_{h \rightarrow \infty}\sin(h)=Undefined\]

OpenStudy (anonymous):

Ok thanks guys that's what i originally thought. My key must be wrong.

OpenStudy (anonymous):

By undefined it is implied that we can't really say what the limit will be because sine will keep bouncing around in the interval [0, 1] as h goes to infinity and won't really "approach" or "converge" to a specific value.

OpenStudy (anonymous):

And lol at your undefended @cwrw238